What minimum work is required to lift a 14 kg mass a distance of 7.0 m? begin mathsize 14px style open parentheses straight g equals 10 space bevelled straight m over straight s squared close parentheses end style
1
Expert's answer
2019-07-01T09:29:40-0400

By the definition of the work done, we have:


"W = Fh = mgh = 14 kg \cdot 7.0 m \cdot 10 \dfrac{m}{s^2} = 980 J."

Answer:

"W = 980 J."